OpenStudy (anonymous):

simplify 5 4/10-3 1/3=

OpenStudy (anonymous):

i get the multiply the 5*10+4 to get the 54/10 & the 3*3+1=10/3 but from there us were i lose it

OpenStudy (anonymous):

it will become......10*5+4/10=54/10 and 3*3+1/3=10/3 now, we have to find the least common multiple so it is 30 now, 54*3+10*10/30=162+100/30=262/30..

OpenStudy (anonymous):

or we can reduce it to....262/30=131/15

OpenStudy (anonymous):

so then i would take the 15 divide the 131 right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

just i divide by numerator and denominator by 2...so we got 131/15

OpenStudy (anonymous):

ok maybe i got it! lol thank you sooooooo much

OpenStudy (anonymous):

You are welcome:)

OpenStudy (anonymous):

u r a lifesaver!